# Distilling heterogeneous treatment effects: Stable subgroup estimation in causal inference

### v1

@@ Appendix E Additional Teacher Model Selection Simulation Results
In this section, we provide the Jaccard SSI simulation results (Figures A9-A14), used to select the teacher model, for each of the subgroup DGPs studied in Section 5 and Appendix D. As discussed in Section 5.3, higher Jaccard SSI generally corresponds to more accurate subgroup estimation regardless of the choice of subgroup DGP.
In practice, we reiterate that substantive researchers should leverage their domain knowledge when choosing the tree depth(s) $d$ to consider in the teacher model selection procedure. In cases where such prior knowledge is limited, we recommend that researchers view the Jaccard SSI results alongside the feature stability distributions and other diagnostic tools (see Appendix A.7) to gain a more holistic perspective. This holistic view of the distilled method and its stability can often shed light on an appropriate choice of tree depth. For example, the feature stability distributions in Figures A9-A14 clearly illuminate that the trees grown to depths 3 and 4 are substantially more unstable than the depth-2 trees. That is, the distribution of features selected at depths 3 and 4 are far more heterogeneous than that at depth 2. As with the Jaccard SSI, a more stable or homogeneous feature distribution is generally a positive sign and may serve as heuristic to help choose the tree depth.

### v2

@@ Appendix E Additional Teacher Model Selection Simulation Results
In this section, we provide the Jaccard SSI simulation results (Figures A10-A15), used to select the teacher model, for each of the subgroup DGPs studied in Section 5 and Appendix D. As discussed in Section 5, higher Jaccard SSI generally corresponds to more accurate subgroup estimation regardless of the choice of subgroup DGP.
In practice, we reiterate that substantive researchers should leverage their domain knowledge when choosing the tree depth(s) $d$ to consider in the teacher model selection procedure. In cases where such prior knowledge is limited, we recommend that researchers view the Jaccard SSI results alongside the feature stability distributions and other diagnostic tools (see Appendix A.8) to gain a more holistic perspective. This holistic view of the distilled method and its stability can often shed light on an appropriate choice of tree depth. For example, the feature stability distributions in Figures A10-A15 clearly illuminate that the trees grown to depths 3 and 4 are substantially more unstable than the depth-2 trees. That is, the distribution of features selected at depths 3 and 4 are far more heterogeneous than that at depth 2. As with the Jaccard SSI, a more stable or homogeneous feature distribution is generally a positive sign and may serve as heuristic to help choose the tree depth.
@@ F.1 Applying Jaccard SSI
F.2 Analyzing the subgroup treatment effects
[…]
Subgroup Results
[…]
F.3 Stability analysis
To consider the stability of the different methods, we perform a bootstrap analysis. Across 100 different bootstrap samples, we apply CDT, virtual twins, as well as tree-based methods and linear methods, to compare the different subgroups obtained under each technique. For each bootstrap instance, we compute the proportion of times features were used to construct subgroups. We summarize the results across these runs in Figure A17.
Figure A17: For our case study, we summarize the proportion of times across the 100 bootstrap instances features are split upon to form rules and compose subgroups across methods (colors). Proportions closer to one indicate more stability and robustness to small data perturbations.
Causal tree also similarly constructs subgroups using CD8 cell count, participants’ starting weight, and the number of days of pre-175 antiretroviral therapy. However, it additionally splits on other covariates, depending on the bootstrap iteration, with less consistency. Similarly, virtual twins consistently constructs subgroups on those same three features, but also constructs substantially more subgroups using other features, resulting in more subgroups that are challenging to interpret.
[…]